OpenStudy (anonymous):

Find the x-coordinates of the points of intersection for the following functions below: f(x) = x^2-4x g(x) = -x^2+4x+3

OpenStudy (anonymous):

i did this and used the quadratic formula

OpenStudy (anonymous):

Set them equal: x^2-4x=-x^2+4x+3 2x^2-8x+3=0

OpenStudy (anonymous):

that should be -3**

OpenStudy (anonymous):

i go tthat

OpenStudy (anonymous):

I then landed up with 8+- sqrt88 all over 4

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

well all the answers have sqrt 22 in them :-(

OpenStudy (anonymous):

Well, because sqrt(88)=2(sqrt(22))

OpenStudy (anonymous):

cause 4 times 22

OpenStudy (anonymous):

Yup :D

OpenStudy (anonymous):

so 2-1/2sqrt22 and then the plus side

OpenStudy (anonymous):

yay!!!